A teachers looks over her class and notices a few trends. Out of the 60 students in the class, 23 have brown hair, seven have green eyes, and three have both brown hair and green eyes. What is the probability that a randomly selected student has either brown hair or green eyes?
Answer #1

P(Either brown hair or green eyes)

= P(brown hair) + P(Green eyes) - P(Both)

= 23/60 + 7/60 - 3/60

= 27/60

= 9/20

= 0.45